Write the slope-intercept equation for the line that passes through
(2,-9) and (4,1)

Answer by mananth(16946)


x1 y1 x2 y2
2 -9 4 1

slope m = (y2-y1)/(x2-x1)
( 1 - -9 )/( 4 - 2
( 10 / 2 )
m= 5

Plug value of the slope and point ( 2,-9) in
Y = m x + b
-9.00 = 10 + b
b= -9 - 10
b= -19
So the equation will be
Y = 5 x -19
m.ananth@hotmail.ca